PHENIX Run 16 Status Meeting Highlights - April 18, 2016

phenix run16 status time meeting 04 18 2016 n.w
1 / 8
Embed
Share

"Discover the latest updates from the PHENIX Run 16 status meeting held on April 18, 2016. Topics covered include luminosity, drift chamber issues, maintenance schedules, and strategies to achieve the required integrated luminosity rate. Stay informed with the key details discussed during the meeting."

  • PHENIX
  • Status Meeting
  • Luminosity
  • Drift Chamber
  • Maintenance

Uploaded on | 0 Views


Download Presentation

Please find below an Image/Link to download the presentation.

The content on the website is provided AS IS for your information and personal use only. It may not be sold, licensed, or shared on other websites without obtaining consent from the author. If you encounter any issues during the download, it is possible that the publisher has removed the file from their server.

You are allowed to download the files provided on this website for personal or commercial use, subject to the condition that they are used lawfully. All files are the property of their respective owners.

The content on the website is provided AS IS for your information and personal use only. It may not be sold, licensed, or shared on other websites without obtaining consent from the author.

E N D

Presentation Transcript


  1. PHENIX Run16 status time meeting 04/18/2016 Denis Jouan PHENIX Run 16 Coordinator Institut de Physique Nucl aire Orsay, CNRS/IN2P3, universit Paris sud, Universit Paris Saclay

  2. luminosity In way to reach sooner the BUP goal 1.8nb-1 narrow vertex 04/18/2016 CAD time meeting: PHENIX status 2

  3. 04/18/2016 CAD time meeting: PHENIX status 3

  4. 04/18/2016 CAD time meeting: PHENIX status 4

  5. Example: last evening Some room for more, but already 4300Hz after 8 hours Quite a beautiful shape 5500Hz after 8 hours -500 Hz/hour There seem to be a sense in this history 04/18/2016 CAD time meeting: PHENIX status 5

  6. Drift chamber WEST SIDE : broken wire problem Some problems with detectors Drift chamber west: wires broken 7april 16 april 20% acceptance lost Would require days 17 april: HV PS on drift chamber east Mainframe changed Sunday morning 04/18/2016 CAD time meeting: PHENIX status 6

  7. Tomorrow morning access Tof E: LV problem Several modules taken out, in process of being repaired (> 1 hour access, from Wednesday morning. +crate problem ? FVTX: needs 3 hours access for noise (dropout) improvement MuTr: a noisy octant from yesterday. Could be connectors. 04/18/2016 CAD time meeting: PHENIX status 7

  8. Summary Integrated luminosity rate above BUP goal needed rate Could accept more luminosity at the end of store. (beta*, 56MHz, ?) (even if at this time of run Keeping full detector efficiency require constant action (regular accesses) 04/18/2016 CAD time meeting: PHENIX status 8

More Related Content